Find the slope-intercept form of the equation of the line passing through the point (2,-1) and 1. perpendicular 2. parellel to the line -4x+2y=7? Thanks in advance.
Answer by checkley71(8403)

-4x+2y=7
2y=4x+7
y=4x/2+7/2
y=2x+7/2 this line has a slope of 2 so the line equation through (2,-1) & a slope of 2 is:
the perpendicular line through (2,-1) has a slope of -1/2 or -.5. thus:
-1=-.5*2+b
-1=-1+b
b=0 thus the equation iis:
y=-.5x for the perpendicular line through (2,-1) & a slope of -.5.
--------------------------------------------------------------------
for the parallel line we have:
-1=2*2+b
-1=4+b
b=-5 so the equation of a parallel line through (2,-1) is:
y=2x-5